Shift report 20030209 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Feb 09 2003 - 07:39:20 EST

  • Next message: Apache: "Shift report 20030209 08:00-16:00"
    ------------------------------------------------------------
    Date: Sun Feb  9 07:39:19 2003, Shift: 00:00-08:00
    Supervisor: PS, with: 
    Log book#14 pages 6 - 6
    
    Shift Summary:
    ------------------------------------------------------------
    
    1. We had a useful beam from 1:30 until
    around 6:00. The beam was supposed to be dumped at 6:00
    but finally, was dumped at 7:00. The delay in dump was due 
    to abort kicker problem. Between 6:00 and 7:00 the beam was useless (TOFW ~150k).
    
    2. Data runs are 7844, 7845, 7856. Summing these runs we get
    430k of trig. 3 and 70k(8deg.) + 47k(12deg.) of trig. 6.
    
    3. No detector problems detected. 
       FS moved to 12 deg.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sun Feb 09 2003 - 07:43:34 EST